40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Apr 30 21:05:32 2012, Jenne, Update, SUS, ETMY oplev power supply dead 
    Reply  Mon Apr 30 21:24:43 2012, Koji, Update, SUS, ETMY oplev power supply dead 
       Reply  Mon Apr 30 22:41:34 2012, Jenne, Update, SUS, ETMY oplev power supply dead 
          Reply  Fri May 18 11:19:17 2012, steve, Update, SUS, ETMY oplev laser is dead 
             Reply  Tue May 22 08:03:01 2012, steve, Update, SUS, ETMY oplev gets new laser 
                Reply  Tue May 29 15:55:22 2012, steve, Update, SUS, ETMY oplev spot size reduced ETMYoplev.jpgIMG_1243.JPG
                   Reply  Tue May 29 21:34:08 2012, Jenne, Update, SUS, ETMY oplev spot size reduced 
                   Reply  Wed Jun 6 14:36:58 2012, steve, Update, SUS, ETMY oplev work finished 
                   Reply  Mon Oct 1 17:03:01 2012, steve, Update, SUS, ETMY oplev relayed IMG_1673oplevETMY.jpg
Message ID: 6588     Entry time: Mon Apr 30 21:24:43 2012     In reply to: 6587     Reply to this: 6589
Author: Koji 
Type: Update 
Category: SUS 
Subject: ETMY oplev power supply dead 

And has the temp controller of the green been recovered too? (Push enable button)

 

ELOG V3.1.3-